ANISOTROPY-DEPENDENT POLARIZATION SPECTROSCOPY USING A TWO-PHOTON PUMP-PROBE TECHNIQUE

\underline{S. B. BAYRAM}^{a}, R. MARHATTA, P. KOIRALA, and M. BRIEL, Miami University, Physics Department, Oxford, OH 45056.

We have experimentally investigated the anisotropy-dependent polarization spectra for the ^{133}Cs 6s^{2}S_{1/2} → 6p^{2}P_{3/2} → 10s^{2}S_{1/2} transition using a two-photon two-color pulsed pump-probe technique. In the investigation, circular and linear polarization degrees were measured to extract the anisotropy-dependent depolarization cross sections in the excited 6p^{2}P_{3/2} level cesium atoms due to collisions with the ground-level argon atoms over the Zeeman coherences. We also studied anisotropy oscillations during the pump-probe overlap time in the excited state cesium due to the coupling of nuclear momentum to the electronic angular momentum. The spectra show strong depolarization of the 6p^{2}P_{3/2} level and the values for the measured depolarization cross sections are in good agreement with experiment and theoretical predictions.

aSupport by Research Corporation under the grant number CC6119C is gratefully acknowledged.